# From square plaquettes to triamond lattices for SU(2) gauge theory

### v1

@@ III Discussion
The error bars on Fig. 3 are statistical only, and self-mitigation has clearly handled the dominant systematic error. The mitigated result at $\tau=1.3$ happens to be furthest from the true ground state, being almost four standard deviations away, but subsequent time steps agree nicely with the true ground state. This demonstrates a valuable feature of the QITE algorithm. Output from one time step is needed as input for the following step, and yet the algorithm recovers quickly from an outlier.
@@
Another interesting glimpse into the symmetries of the triamond lattice comes from noticing that the six vectors comprising a triamond lattice form a regular tetrahedron. Begin with one gauge link of each color (recall Eqs. (11) and (12)) and translate each one spatially without any rotations. The red, green and blue links will form the equilateral triangular base of the tetrahedron. The cyan, magenta and yellow links rise from the corners of that base to meet at the top of the tetrahedron. As might be expected from the sentences preceding Eq. (25), the cyan link of the tetrahedron touches all links except the red one, the magneta link touches all but green, and the blue link touches all but yellow.
@@
Figure 7: The circuit used to obtain the bottom panel of Fig. 6. The 12 qubits are labeled by their colors, with two copies of each color appearing in the list. A constant angle is defined by $\phi=\arccos(1/\sqrt{3})$.
Figure 8: The circuit used to obtain the center panel of Fig. 6. The 12 qubits are labeled by their colors, with two copies of each color appearing in the list.
The present study has explored error mitigation and the triamond lattice as separate steps toward a common goal. Future work can combine error mitigation with the triamond lattice to continue the progress toward a practical implementation of lattice gauge theories on quantum computers.
IV Methods
IV.1 Tomography for the QITE algorithm
@@
$\left|\Delta_{0}\right>\equiv e^{-i\tau\hat{A}}\left|\psi\right>$ (26)
@@
$\left|\Delta\right>\equiv-i\tau\hat{A}\left|\psi\right>\,.$ (27)
@@
The commutators defining the elements of $b$ use the Hamiltonian from Eq. (1). The numerical computation of $b$, $S+S^{T}$ and $r^{\prime}/r$ (from Eq. (4)) requires nine expectation values. Three of them can be obtained by preparing the state $\left|\psi\right>$ in a quantum computer and measuring each qubit, which gives
@@
IV.2 Randomized compiling for CNOT gates
@@
Table 1: The 16 options for randomized compiling.
Applied before the CNOT Applied after the CNOT
control target control target
I I I I
I X I X
I Y Z Y
I Z Z Z
X I X X
X X X I
X Y Y Z
X Z Y Y
Y I Y X
Y X Y I
Y Y X Z
Y Z X Y
Z I Z I
Z X Z X
Z Y I Y
Z Z I Z
IV.3 Deriving the triamond SU(2) Hamiltonian
@@
$H_{B}=\frac{V}{g^{2}a}\sum_{\vec{w}={\rm white}}{\rm Tr}\big{(}F_{xy}^{2}(\vec{w})+F_{yz}^{2}(\vec{w})+F_{zx}^{2}(\vec{w})\big{)}+O(a)\,,$ (62)
where $V$ is the volume per white site. There are two white sites in each unit cell, and the volume of a unit cell is $(2\sqrt{2}a)^{3}$, so $V=8\sqrt{2}a^{3}$. Therefore the magnetic Hamiltonian becomes
$H_{B}=\frac{\sqrt{2}}{g^{2}a}\sum_{\vec{w}={\rm white}}\left(12-\sum_{k=1}^{6}\mathcal{P}_{k}(\vec{w})\right)+O(a)\,.$ (63)
@@
$H_{E}$ $=$ $\frac{V}{6g^{2}}\sum_{n={\rm links}}{\rm Tr}\left(E_{x}^{2}(n)+E_{y}^{2}(n)+E_{z}^{2}(n)\right)+O(a)$
$=$ $\frac{4\sqrt{2}a^{3}}{3g^{2}}\sum_{n={\rm links}}{\rm Tr}\left(E_{x}^{2}(n)+E_{y}^{2}(n)+E_{z}^{2}(n)\right)+O(a)$
@@
IV.4 Hamiltonian and circuits for triamond unit cell
To obtain the results in Fig. 6, we begin with the Hamiltonian of SU(2) gauge theory truncated to $j\in\{0,\frac{1}{2}\}$ on a unit cell of the triamond lattice with periodic boundary conditions. The general expression was obtained in Sec. II.2 and here it is written in terms of Pauli operators.
@@ IV.4 Hamiltonian and circuits for triamond unit cell
$-\frac{3\sqrt{2}}{4g^{4}a}\left(\prod_{k\in\{r,c\}}\frac{1-Z_{k}}{2}\right)\prod_{k\notin\{r,c\}}X_{k}$
$+\frac{3\sqrt{2}}{16g^{4}a}\left(\prod_{k\in\{r,c\}}\frac{1+Z_{k}}{2}\right)\prod_{k\notin\{r,c\}}X_{k}\,,$
@@
$-\frac{3\sqrt{2}}{4g^{4}a}\left(\prod_{k\in\{g,m\}}\frac{1-Z_{k}}{2}\right)\prod_{k\notin\{g,m\}}X_{k}$
$+\frac{3\sqrt{2}}{16g^{4}a}\left(\prod_{k\in\{g,m\}}\frac{1+Z_{k}}{2}\right)\prod_{k\notin\{g,m\}}X_{k}\,,$
and application of the non-blue non-yellow plaquette.
@@
$-\frac{3\sqrt{2}}{4g^{4}a}\left(\prod_{k\in\{b,y\}}\frac{1-Z_{k}}{2}\right)\prod_{k\notin\{b,y\}}X_{k}$
$+\frac{3\sqrt{2}}{16g^{4}a}\left(\prod_{k\in\{b,y\}}\frac{1+Z_{k}}{2}\right)\prod_{k\notin\{b,y\}}X_{k}\,.$
@@
The next thing needed to obtain the results in Fig. 6 is a quantum circuit to prepare the ansatz containing the variational parameter. Based on the discussion surrounding Eq. 25, we anticipate that the physical ground state will have a contribution from the bare ground state as well as a triply degenerate contribution from three plaquette terms. Our variational parameter $\theta$ will be used to adjust the relative strengths of these two contributions. Our circuit is displayed in Fig. 7 and is responsible for the bottom panel of Fig. 6. This particular ansatz is quite simple, with two qubits at its core and extra CNOT gates adjusting the other qubits in a straightforward way. The circuit has been designed such that it requires minimal connectivity among the 12 qubits and is suitable, for example, to run on IBM’s heavy hex architecture without additional swap gates. The center panel of Fig. 6 uses the trial state obtained from the circuit in Fig. 8. The top panel of Fig. 6 uses a trial state that differs from Fig. 8 simply by appending a Pauli $X$ gate at the end of the circuit for each cyan, magenta and yellow qubit.
@@
Acknowledgements.
The authors thank R. M. Woloshyn for suggesting that self-mitigation be applied to QITE, W. G. Parrott for discussions about quantum imaginary time evolution, and M. Tsang for discussions about triamond crystals. We thank all three for reading a draft of the manuscript. We also acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team. This work is affiliated with the QC4HEP Working Group and supported by Canada’s Natural Sciences and Engineering Research Council.

### v2

@@ Discussion
The error bars in Fig. 3 are statistical only, and self-mitigation has clearly handled the dominant systematic error. The mitigated result at $\tau=1.3$ happens to be furthest from the true ground state, being several standard deviations away, but subsequent time steps agree nicely with the true ground state. This demonstrates a valuable feature of the QITE algorithm. Output from one time step is needed as input for the following step, and yet the algorithm recovers quickly from an outlier.
@@
Another interesting glimpse into the symmetries of the triamond lattice comes from noticing that the six vectors comprising a triamond lattice form a regular tetrahedron. Begin with one gauge link of each color (recall Eqs. (11) and (12)) and translate each one spatially without any rotations. The red, green and blue links will form the equilateral triangular base of the tetrahedron. The cyan, magenta and yellow links rise from the corners of that base to meet at the top of the tetrahedron. As might be expected from the sentences preceding Eq. (25), the cyan link of the tetrahedron touches all links except the red one, the magenta link touches all but green, and the yellow link touches all but blue.
@@
Figure 7: The circuit used to obtain the ground state eigenvalue. The 12 qubits are labeled by their colors, with two copies of each color appearing in the list. A constant angle is defined by $\phi=\arccos(1/\sqrt{3})$. The ground state eigenvalue for one unit cell of the triamond lattice is displayed in Fig. 5(c).
Figure 8: The circuit used to obtain the first excited state eigenvalue. The 12 qubits are labeled by their colors, with two copies of each color appearing in the list. The first excited state eigenvalue for one unit cell of the triamond lattice is displayed in Fig. 5(b).
Fig. 6 shows the successful application of imaginary time evolution to a lattice in three spatial dimensions. The result is a robust preparation of the ground state. The triamond framework offers a systematic path toward larger lattices and relaxed truncations of the gauge theory. Future work in this direction can continue the progress toward a practical implementation of lattice gauge theories on quantum computers.
Methods
Tomography for the QITE algorithm
@@ Methods
$\left|\Delta_{0}\right>\equiv\left(\frac{e^{-i\tau\hat{A}}-1}{\tau}\right)\left|\psi\right>$ (27)
@@
$\left|\Delta\right>\equiv-i\hat{A}\left|\psi\right>\,.$ (28)
@@
The commutators defining the elements of $b$ use the Hamiltonian from Eq. (1) for square plaquettes or Eq. (25) for the triamond unit cell. The numerical computation of $b$, $S+S^{T}$ and $r^{\prime}/r$ (from Eq. (4)) requires nine expectation values. Three of them can be obtained by preparing the state $\left|\psi\right>$ in a quantum computer and measuring each qubit, which gives
@@
Randomized compiling for CNOT gates
@@
Deriving the triamond SU(2) Hamiltonian
@@
$H_{B}=\frac{V}{g^{2}}\sum_{\vec{w}={\rm white}}{\rm Tr}\big{(}F_{xy}^{2}(\vec{w})+F_{yz}^{2}(\vec{w})+F_{zx}^{2}(\vec{w})\big{)}+O(a)\,,$ (63)
where $V$ is the volume per white site. There are two white sites in each unit cell, and the volume of a unit cell is $(2\sqrt{2}a)^{3}$, so $V=16\sqrt{2}a^{3}$. Therefore the magnetic Hamiltonian becomes
$H_{B}=\frac{2\sqrt{2}}{g^{2}a}\sum_{\vec{w}={\rm white}}\left(12-\sum_{k=1}^{6}\mathcal{P}_{k}(\vec{w})\right)+O(a)\,.$ (64)
@@
$H_{E}$ $=$ $\frac{Vg^{2}}{6}\sum_{n={\rm links}}{\rm Tr}\left(E_{x}^{2}(n)+E_{y}^{2}(n)+E_{z}^{2}(n)\right)+O(a)$ (66)
$=$ $\frac{8\sqrt{2}a^{3}g^{2}}{3}\sum_{n={\rm links}}{\rm Tr}\left(E_{x}^{2}(n)+E_{y}^{2}(n)+E_{z}^{2}(n)\right)$
$+O(a)$
@@
Hamiltonian and circuits for triamond unit cell
To obtain the results in Fig. 5, we begin with the Hamiltonian of SU(2) gauge theory truncated to $j\in\{0,\frac{1}{2}\}$ on a unit cell of the triamond lattice with periodic boundary conditions. The general expression was obtained in Results and here it is written in terms of Pauli operators.
@@
$-\frac{3\sqrt{2}}{4g^{4}a}\left(\prod_{k\in\{r,c\}}\frac{1+Z_{k}}{2}\right)\prod_{k\notin\{r,c\}}X_{k}$
$+\frac{3\sqrt{2}}{16g^{4}a}\left(\prod_{k\in\{r,c\}}\frac{1-Z_{k}}{2}\right)\prod_{k\notin\{r,c\}}X_{k}\,,$
@@
$-\frac{3\sqrt{2}}{4g^{4}a}\left(\prod_{k\in\{g,m\}}\frac{1+Z_{k}}{2}\right)\prod_{k\notin\{g,m\}}X_{k}$
$+\frac{3\sqrt{2}}{16g^{4}a}\left(\prod_{k\in\{g,m\}}\frac{1-Z_{k}}{2}\right)\prod_{k\notin\{g,m\}}X_{k}\,,~{}~{}$
and application of the non-blue non-yellow plaquette,
@@
$-\frac{3\sqrt{2}}{4g^{4}a}\left(\prod_{k\in\{b,y\}}\frac{1+Z_{k}}{2}\right)\prod_{k\notin\{b,y\}}X_{k}$
$+\frac{3\sqrt{2}}{16g^{4}a}\left(\prod_{k\in\{b,y\}}\frac{1-Z_{k}}{2}\right)\prod_{k\notin\{b,y\}}X_{k}\,.$
@@
The next thing needed to obtain the results in Fig. 5 is a quantum circuit to prepare the ansatz containing the variational parameter. Based on the discussion surrounding Eq. (25), we anticipate that the physical ground state will have a contribution from the bare ground state as well as a triply degenerate contribution from three plaquette terms. Our variational parameter $\theta$ will be used to adjust the relative strengths of these two contributions. Our circuit is displayed in Fig. 7 and is responsible for the bottom panel of Fig. 5. This particular ansatz is quite simple, with two qubits at its core and extra CNOT gates adjusting the other qubits in a straightforward way. The circuit has been designed such that it requires minimal connectivity among the 12 qubits and is suitable, for example, to run on IBM’s heavy hex architecture without additional swap gates. The center panel of Fig. 5 uses the trial state obtained from the circuit in Fig. 8. The top panel of Fig. 5 uses a trial state that differs from Fig. 8 simply by appending a Pauli $X$ gate at the end of the circuit for each cyan, magenta and yellow qubit.
@@
Data availability
The numerical data for generating the figures are available from the authors upon request.
Code availability
Code for calculating self-mitigated imaginary time evolution is provided at https://github.com/randylewis/SelfMitigatedQITE.